# Chain-of-Thought Reasoning Improves Context-Aware Translation with Large Language Models

### v1

@@ 5.1.  Prompts: Reasoning or Not
Prompts to LLMs are typically made of a system prompt, which indicates the role, persona, or style expected for the answer, and a user prompt, which specifies the task, ending with data for the task. Here, we instruct the LLMs to solve the contrastive task, i.e. select the correct translation, with several options for the system and user prompts.
@@
The system prompt can be either empty, simple, or detailed. The simple version is shown in Table 2, while the detailed version is twice longer and contains more constraints, but no instruction on how to solve the task. These prompts are the same for the two benchmarks, anaphora or lexical cohesion.
@@
The user prompt can be either a simple definition of the task, which includes the data as variables that are instantiated for each test item, or instructions on how to reason step-by-step to solve it. From lack of space, we only show in Table 3 the prompt with the step-by-step reasoning instructions for the anaphora task. The reasoning prompt for the lexical cohesion task differs from the anaphora one at the third step: “Step 3 - Find the text in English line 1 which is identical to the text found at Step 2”. [^7] These instructions are oriented towards solving the anaphora or lexical cohesion problem, and are therefore more focused than the document-level step-by-step translation considered by Briakou et al. (2024).
When providing the answer, strictly follow this format:
'''Choice: (1 or 2)'''
Table 2: System prompt: simple version.
Step 1 - Find the difference between the two translations of line 2.
Step 2 - Find the text in English line 2 which is the cause of the translation difference at Step 1.
Step 3 - Find the text in English line 1 to which the text found at Step 2 refers.
Step 4 - Find how the text from Step 3 is translated in French translation of English line 1.
Step 5 - Find the correct word in French to refer to this text.
Step 6 - Find which French translation includes this word.
Select ’1’ if French translation number 1 is more correct, or answer ’2’ if French translation number 2 is more correct.
Table 3: User prompt for the anaphora task: step-by-step reasoning instructions.
@@
Figure 1: Mean accuracy (top bars) and inconsistency (bottom bars, smaller are better) on the contrastive anaphora task. For each LLM, the four bars are those of the prompts: (none, simple), (simple, simple), (detailed, simple) and (simple, step-by-step). LLMs are ranked by decreasing scores with the last prompt (yellow), which reaches highest overall performance.
5.2.  Results
@@
We designed the prompts through a series of experiments on half of the benchmark data (development set) with 12 LLMs. We settled on four combinations of system and user prompts: (none, simple), (simple, simple), (detailed, simple) and (simple, step-by-step). The accuracy and inconsistency scores of these combinations are shown in Figure 1 on the anaphora task for all 12 LLMs. The numbers are given in Table 4 for the 5 most interesting LLMs, on the development and test sets. The systems are ranked by decreasing accuracy of the best overall prompt, which is the (simple, step-by-step) one, i.e. when models are encouraged to reason explicitly before answering. The findings from experiments with the validation sets are confirmed on the test sets. From lack of space, we do not provide the graph for the lexical task, but only show scores for 5 LLMs on both parts of the benchmark data in Table 5.
@@
Table 4: Comparison of accuracy and inconsistency on development and test sets for the anaphora task. For each model, the four lines of scores are those of the prompts: (none, simple), (simple, simple), (detailed, simple) and (simple, step-by-step).
@@
Table 5: Comparison of accuracy and inconsistency on validation and test sets for the lexical choice task. For each model, the four lines of scores are those of the prompts: (none, simple), (simple, simple), (detailed, simple) and (simple, step-by-step).
5.3.  Findings of the Contrastive Task
To the best of our knowledge, the performance of previous systems on this task is as follows. The creators of DiscEvalMT designed a system which encodes and decodes jointly the pairs of sentences, reaching accuracies of 0.72 and 0.57 respectively on anaphora and lexical cohesion (Bawden et al., 2018, Table 2). In a comparative evaluation study, Lopes et al. (2020, Table 7) found that the best performing approach was the concatenation of both sentences on both sides, as proposed by Tiedemann and Scherrer (2017), reaching 0.82 and 0.55 accuracies. The encoder-decoder model proposed by Pal et al. (2024, Table 4), which used a separate encoder for the context, reached 0.54 and 0.52 accuracy, while Zhang et al. (2022, Table 5) announced 0.64 accuracy on the anaphora task. A quality estimation approach, proposing a contextual version of the COMET-QE metric (Vernikos et al., 2022), reached respectively 0.83 and 0.68 accuracy.
On the anaphora task (Table 4), the best results on the validation set came from GPT-4o with the reasoning prompt and from GPT-4 with the simplest prompt (no reasoning), both reaching 0.88 accuracy. However, GPT-4o with reasoning outperformed GPT-4 on the test set, reaching a nearly perfect accuracy of 0.97. For both systems, the presentation order of the alternatives had almost no influence on the answer, with an inconsistency of only 0.1. GPT-4 also did well with the other prompts, except the reasoning one, scoring up to 0.86 with a consistency of 0.04, and GPT-4-turbo followed closely. GPT-3.5-turbo showed a larger variability, with inconsistency scores up to 0.45 on the validation set.
Among LLMs run locally, Phi-4 stood out with 0.82 accuracy with the reasoning prompt and very stable outputs (inconsistency of 0.01), making it the most reliable open-source model. DeepSeek-R1 32B peaked at 0.74 accuracy also with a low inconsistency of 0.02, but smaller DeepSeek variants like the 14B and 8B models dropped to around 0.60 and were less consistent.
@@ 5.3.  Findings of the Contrastive Task
Overall, higher accuracy goes hand-in-hand with better consistency. The reasoning prompt was helpful for models that could handle it. For smaller or less instructed models, more complex prompts often led to worse outcomes or to output formatting issues.
On the lexical choice task (Table 5) the trends are similar. The scores are even higher than for anaphora, with excellent consistency. The best performance came from GPT-4o with the reasoning-style prompt, reaching an accuracy of 0.96 and a very low inconsistency of 0.02. GPT-4 also performed strongly, with up to 0.92 accuracy, although its results varied slightly more across prompts. Again, models with strong reasoning capabilities, such as GPT-4o and GPT-4, consistently outperformed others, especially when the prompt encouraged CoT reasoning, and among open-source options, Phi-4 offers the best combination of performance and efficiency.
In terms of the length of response (a correlate of cost for OpenAI models), DeepSeek-R1 models consistently generated long responses (over 600 tokens), regardless of the prompt, as they are trained to perform CoT reasoning. In contrast, most other models had output lengths that were more clearly shaped by the prompt. LLaMA 3.3 and Phi-4 produced some of the longest reasoning responses apart from DeepSeek-R1, with 279 and 227 tokens on average under the reasoning prompt. On the contrary, GPT-3.5-turbo, Mistral, and LLaMA 3.2 generated very short outputs, even under reasoning prompts, rarely exceeding 10 tokens. Therefore, given the higher cost of CoT responses, reasoning prompts are only worth using when they improve performance, as in the case of GPT-4o and Phi-4.
@@ 6.  Translation Task
In this section, we test the ability of LLMs to correctly translate a sentence when inter-sentential constraints, i.e., pronominal anaphora and lexical cohesion, are involved. Indeed, as observed by Post and Junczys-Dowmunt (2024) for encoder-decoder MT systems, scores on the contrastive task are not necessarily correlated to translation quality.
Using the same dataset, we give to each LLM the two source sentences (EN) and the reference translation of the first sentence (FR). Using several possible prompts, we obtain the translation of the second sentence from the LLM and score it against the reference translation present in the dataset using the four evaluation metrics presented in Section 3.2 above: BLEU, chrF, BERTScore, and COMET. As prompt engineering relied less on the development set as than for the first task, and we report below the results on the full set of 200 examples for each task.
@@ 6.1.  Prompts: Reasoning or Not
Our goals are again to determine the best performance of LLMs on the two benchmarks, and to find if CoT reasoning improves performance over translation with no reasoning. The no-reasoning prompt is the same for both benchmarks: the system part is shown in Table 6, while the user part simply provides the two EN sentences and the FR translation of the first one.
### Instructions:
- You will also be given the French translation of the context sentence.
- Translate only the second English sentence into French.
- Return only the French translation of the second sentence.
- Do not include any explanation or additional text.
Table 6: System prompt without reasoning for the translation task.
For the reasoning prompt, we state in the system prompt several steps that guide the reasoning for each task, shown for both benchmarks in Table 7. An additional instruction at the end (not shown here) requires the LLM to enclose the reasoning between XML-like tags. We found that this instruction led to similar or better performance and an easier to parse output in comparison to an unstructured one.
[Same as Table 6 except its last two lines.]
1. Identify pronouns/references in the second sentence.
2. Find their referent in the first sentence.
3. Check how that referent is translated in the French context.
4. Choose the correct French pronoun/reference.
[Same as Table 6 except its last two lines.]
To ensure lexical consistency:
1. Identify any key word or expression in the second sentence in English that could be translated in more than one way into French.
2. Check if the translation of the first sentence into French already provides a preferred translation for that key word or expression.
Use the same choice if appropriate for your translation of the second sentence into French.
Table 7: System prompts with reasoning instructions for the translation task on the anaphora benchmark (top) and the lexical cohesion one (bottom).
6.2.  Results
The translation scores on the anaphora benchmark, with or without reasoning, and their difference ($\Delta$), are shown in Table 9. For the lexical choice benchmark, the ranking of the LLMs is similar; from lack of space, we only show the scores of the two best systems in Table 10. These numbers are discussed in the next section.
BLEU chrF BERT COMET
Score
Anaphora $r_{P}$ 0.81 0.76 0.60 0.59
$r_{S}$ 0.68 0.72 0.64 0.74
Lexical $r_{P}$ 0.40 0.21 0.31 0.30
choice $r_{S}$ 0.52 0.27 0.30 0.25
Table 8: Pearson ($r_{P}$) and Spearman ($r_{S}$) correlations between baseline scores (without reasoning) and improvements due to reasoning ($\Delta$).
@@ 6.2.  Results
DS-R1 8B 24 21 -2.48 50 45 -4.75 .86 .77 -0.0908 .80 .77 -0.0298
Table 9: Translation quality scores on the anaphora benchmark: without reasoning (w/o), with reasoning (w/), and difference ($\Delta$) between the latter and the former. Values for w/o and w/ are rounded to 2 digits, and $\Delta$ values to 4. Positive values of $\Delta$ indicate progress due to reasoning. ‘DS’ stands for DeepSeek.
@@
Table 10: Translation quality scores (rounded) on the lexical choice benchmark for the two best LLMs.
On both the anaphora and lexical choice experiments, we observe a surprising effect. Unlike many techniques which tend to improve weaker models but do not benefit the top-scoring ones, here reasoning is more beneficial to the LLMs which already score highest without it. Table 8 shows the Pearson correlations and the Spearman rank correlations between baseline scores without reasoning and the improvements obtained with reasoning ($\Delta$), for both tasks. For anaphora, high coefficients for all metrics (0.59–0.81) confirm a consistent “wise-get-wiser” pattern: models that start from higher baselines tend to benefit more from reasoning. The effect is weaker for lexical choice, although correlations remain positive (0.21–0.52).
Moreover, we confirm that the variations of translation scores between the prompts without reasoning and those with reasoning are consistent across all metrics. This is important because the scoring of LLM translations relies on MT metrics which correlate imperfectly with human preferences, unlike the accuracy metric of the first part. Figure 2 shows correlations among the $\Delta$ values of all systems, between all pairs of metrics. The large coefficients (Pearson: 0.58–0.91, Spearman: 0.66–0.92) indicate that when a model improves (or degrades) on one metric, it usually improves (or degrades) on others as well. Similar values are observed for the lexical choice task.
Figure 2: Correlations between $\Delta$ values (reasoning minus no_reasoning) obtained with various MT metrics for 12 LLMs, on anaphora translation. Each cell shows Pearson (upper part) and Spearman (lower part) correlations.
6.3.  Findings
@@
The results in Table 9 reveal a clear pattern in how models perform on the anaphora benchmark. The GPT family delivers the strongest performance, and the largest increase when reasoning is encouraged. The smaller “turbo” versions benefit most from reasoning in terms of BERTScore and COMET. Phi-4 exhibits the highest improvement on all metrics when reasoning is used. Together with Llama 3.3, they are the highest scoring open-weight models, though Llama 3.3 draws only modest benefits from reasoning. The DeepSeek models present a mixed picture: only the largest model improves its BLEU score, though not the other scores. Models with weaker baselines i.e., DeepSeek-R1 8B and 14B, LLaMA 3.1 and 3.2, and Mistral decline across all metrics when prompted to reason.
@@
For lexical choice, the results in Table 10 reveal a very similar picture. Within the GPT family, reasoning improves GPT-4, GPT-4o, and GPT-3.5-turbo across most metrics. While GPT-4-turbo shows an unexpected small decline, GPT-4 seems to make a slightly more effective use of the reasoning instructions, despite GPT-4o’s stronger baseline score. Among open-weight models, Phi-4 again improves on all metrics with the CoT prompt, while LLaMA 3.3 and DeepSeek-R1 32B improve only in BLEU and chrF, Mistral gains slightly, and the other models degrade. Lexical choice may pose a more difficult challenge compared to pronominal anaphora, as possible translations of ambiguous words are more numerous. This may explain why improvements through reasoning are slightly less consistent for the lexical choice task.

### v2

@@ 5.1.  Prompts: Reasoning or Not
Prompts to LLMs are typically made of a system prompt, which indicates the role, persona, or style expected for the answer, and a user prompt, which specifies the task, ending with data for the task. Here, we instruct the LLMs to solve the contrastive task, i.e. select the correct translation, with several variants for the system and user prompts.
@@
The system prompt can be either empty (no system prompt), simple, or detailed. Each of these versions is shown in Appendix A.1. The detailed version is twice longer than the simple one and contains more constraints, but neither of them gives any instruction on how to solve the task. The system prompts are the same for the two benchmarks, anaphora or lexical cohesion, on the contrastive task.
@@
The user prompt can be either a simple definition of the task, which includes the data, or instructions on how to reason step-by-step to solve it. These prompts are also shown in Appendix A.1. The step-by-step reasoning instructions for the anaphora and the lexical choice tasks are very similar: the main difference is at the third step. For anaphora, this is “Find the text in English line 1 to which the text found at Step 2 refers” while for lexical cohesion this is “Find the text in English line 1 which is identical to the text found at Step 2”. [^8] These instructions are oriented towards solving the anaphora or lexical cohesion problems, and are therefore more focused than the document-level step-by-step translation considered by Briakou et al. (2024).
We designed the prompts through a series of experiments on half of the benchmark data (the development set) with 12 LLMs. We settled on four combinations of system and user prompts: (1) no system prompt and simple user prompt; (2) simple system and user prompts; (3) detailed system prompt and simple user prompt; (4) simple system prompt and step-by-step (reasoning) user prompt.
The DeepSeek-R1 models are “reasoning” models, as they were trained to generate a detailed solution with multiple verifications and to conclude with the final answer. The usage instructions for DeepSeek-R1, which we follow, recommend to avoid a system prompt, but to include the words “please reason step by step” and indicate the explicit markup of the answer in the user prompt.
@@
Figure 1: Mean accuracy (top bars) and inconsistency (bottom bars, smaller are better) on the contrastive anaphora task, validation set. For each LLM, the four bars are those of the prompts: (1) no system prompt and simple user prompt; (2) simple system and user prompts; (3) detailed system prompt and simple user prompt; (4) simple system prompt and step-by-step (reasoning) user prompt. LLMs are ranked by decreasing scores obtained with the last prompt (yellow), which reaches highest overall performance.
Figure 2: Mean accuracy (top bars) and inconsistency (bottom bars, smaller are better) across prompt configurations for each model on the lexical choice validation set.
5.2.  Results of the Contrastive Task
@@
The accuracy and inconsistency scores of these combinations of prompts are shown in Figure 1 on the anaphora task, for all 12 LLMs. The numbers are given in Table 2 for the five most interesting LLMs, on the development and test sets. The systems are ranked by decreasing accuracy of the best overall prompt, which is the one with simple system prompt and reasoning user prompt. The findings from experiments with the validation sets are confirmed on the test sets. Similarly, the scores for the lexical choice task are shown in Figure 2, and the scores for five LLMs on both subsets of the benchmark data are shown in Table 3.
@@
Table 2: Comparison of accuracy and inconsistency on development and test sets for the anaphora task. For each LLM, the four lines of scores are those of the prompts: (1) no system prompt and simple user prompt; (2) simple system and user prompts; (3) detailed system prompt and simple user prompt; (4) simple system prompt and step-by-step (reasoning) user prompt.
@@
Table 3: Comparison of accuracy and inconsistency on validation and test sets for the lexical choice task. For each LLM, the four lines correspond to the four configurations of prompts.
To the best of our knowledge, the performance of previous systems is as follows. The creators of DiscEvalMT designed a system which encodes and decodes jointly the pairs of sentences, reaching accuracies of 0.72 and 0.57 respectively on anaphora and lexical cohesion (Bawden et al., 2018, Table 2). In a comparative evaluation study, Lopes et al. (2020, Table 7) found that the best performing approach was the concatenation of both sentences on both sides, as proposed by Tiedemann and Scherrer (2017), reaching 0.82 and 0.55 accuracies. The encoder-decoder model proposed by Pal et al. (2024, Table 4), which used a separate encoder for the context, reached 0.54 and 0.52 accuracy, while Zhang et al. (2022, Table 5) announced 0.64 accuracy on the anaphora task. A quality estimation approach with the COMET-QE metric (Vernikos et al., 2022) reached respectively 0.83 and 0.68 accuracy. Finally, our own use of NLLB-200 only reached, respectively, 0.66 and 0.61 accuracy (first line of Table 2).
On the anaphora task (Figure 1 and Table 2), the best results on the validation set came from GPT-4o with the reasoning prompt and from GPT-4 with the simplest prompt (no reasoning), both reaching 0.88 accuracy. However, GPT-4o with reasoning outperformed GPT-4 on the test set, reaching a nearly perfect accuracy of 0.97. For both systems, the presentation order of the alternatives had almost no influence on the answer, with an inconsistency score of only 0.1. GPT-4 also did well with the other prompts, except the reasoning one, scoring up to 0.86 with an inconsistency of 0.04, and GPT-4-turbo followed closely. GPT-3.5-turbo showed a larger variability, with inconsistency scores up to 0.45 on the validation set.
Among LLMs run locally, Phi-4 stood out with 0.82 accuracy with the reasoning prompt and very stable outputs (inconsistency of 0.01), making it the most reliable open-source model. DeepSeek-R1 32B peaked at 0.74 accuracy also with a low inconsistency of 0.02, but smaller DeepSeek variants like the 14B and 8B models dropped to around 0.60 and were less consistent. Because Deepseek-R1 are reasoning models, the reasoning prompt did not bring significant benefits, apart from a small increase likely related to the instruction to “reason step-by-step” recommended by its authors. As expected from the usage instructions, a detailed system prompt was detrimental.
@@
Overall, higher accuracy goes hand-in-hand with better consistency. The reasoning prompt was helpful for models that could handle it. For smaller or less instructed models, more complex prompts often led to worse outcomes or formatting issues.
On the lexical choice task (Figure 2 and Table 3) the trends are similar. The scores are even higher than for anaphora, with excellent consistency. The best performance came from GPT-4o with the reasoning-style prompt, reaching an accuracy of 0.96 and a very low inconsistency of 0.02. GPT-4 also performed strongly, with up to 0.92 accuracy, although its results varied slightly more across prompts. Again, models with strong reasoning capabilities, such as GPT-4o and GPT-4, consistently outperformed others, especially with the CoT prompt, and among open-source options, Phi-4 offered the best combination of performance and efficiency.
In terms of the length of response, which correlates with cost, DeepSeek-R1 models consistently generated long responses (over 600 tokens), regardless of the prompt, as they are trained to perform CoT reasoning. In contrast, most other models had output lengths that were more clearly shaped by the prompt. Llama 3.3 and Phi-4 produced some of the longest reasoning responses apart from DeepSeek-R1, with 279 and 227 tokens on average under the reasoning prompt. On the contrary, GPT-3.5-turbo, Mistral, and Llama 3.2 generated very short outputs, even with reasoning prompts, rarely exceeding 10 tokens. Therefore, given the higher cost of CoT responses, reasoning prompts are only worth using when they improve performance, as in the case of GPT-4o and Phi-4.
@@ 6.  Translation Task
We now test the ability of LLMs to correctly translate a sentence when inter-sentential constraints – pronominal anaphora and lexical cohesion – are involved. Indeed, as observed by Post and Junczys-Dowmunt (2024) for encoder-decoder MT systems, scores on the contrastive task are not necessarily correlated with translation quality.
Using the same dataset, we give each LLM the two source sentences (EN) and the reference translation of the first sentence (FR). We obtain the translation of the second sentence from the LLM and score it against the reference translation present in the dataset using the four evaluation metrics presented in Section 3.2 above: BLEU, chrF, BERTScore, and COMET. As prompt engineering relied less on the development set than for the first task, we report below the results on the full set of 200 examples for each task.
@@ 6.1.  Prompts: Reasoning or Not
Our goals are again to determine the best performance of LLMs on the two benchmarks, and to find if CoT reasoning improves performance over translation with no reasoning. The no-reasoning prompt is the same for both benchmarks, as shown in Appendix B.1. The system part contains general instructions for translation, while the user part simply provides the two EN sentences and the FR translation of the first one.
For the reasoning prompts (also in Appendix B.1), we make explicit in the system prompt several steps that guide the reasoning for each task. An additional instruction at the end requires the LLM to enclose the reasoning between XML-like tags, which leads to an easier to parse output, as well as similar or better performance.
@@
DeepSeek-R1 8B 24 21 -2.48 50 45 -4.75 .86 .77 -0.0908 .80 .77 -0.0298
Table 4: Translation quality scores on the anaphora benchmark: without reasoning (w/o), with reasoning (w/), and difference ($\Delta$) between the latter and the former. Values for w/o and w/ are rounded to 2 digits, and $\Delta$ values to 4. Positive values of $\Delta$ indicate progress due to reasoning.
@@
gpt-4-turbo 50 49 -0.31 66 65 -0.83 .91 .89 -0.0209 .87 .86 -0.0126
gpt-3.5-turbo 47 49 +2.11 65 66 +0.81 .91 .91 +0.0041 .86 .86 -0.0059
Llama3.3 46 47 +0.15 62 63 +0.90 .90 .90 -0.0018 .85 .84 -0.0030
Phi-4 43 44 +1.77 60 61 +1.74 .89 .90 +0.0064 .83 .84 +0.0021
DeepSeek-R1 32B 39 41 +2.07 56 58 +2.01 .88 .88 -0.0011 .82 .81 -0.0054
DeepSeek-R1 14B 38 33 -5.10 55 51 -3.99 .87 .85 -0.0226 .80 .77 -0.0330
Llama3.1 35 29 -5.71 52 49 -3.53 .87 .86 -0.0144 .79 .77 -0.0261
DeepSeek-R1 8B 31 30 -0.63 48 48 -0.28 .86 .83 -0.0227 .76 .75 -0.0115
Mistral 27 28 +1.17 45 45 +0.11 .84 .85 +0.0046 .74 .75 +0.0101
Llama3.2 25 22 -2.81 42 43 +0.63 .83 .82 -0.0121 .73 .70 -0.0256
Table 5: Translation quality scores on the lexical choice benchmark: without reasoning (w/o), with reasoning (w/), and difference ($\Delta$) between the latter and the former. Values for w/o and w/ are rounded to 2 digits, and $\Delta$ values to 4. Positive values of $\Delta$ indicate progress due to reasoning.
6.2.  Results of the Translation Task
@@
On the anaphora benchmark, the translation scores with or without reasoning, and their differences ($\Delta$), are shown in Table 4. The results show that the GPT family delivers the strongest performance, and the largest increase when reasoning is encouraged. The smaller “turbo” versions benefit most from reasoning in terms of BERTScore and COMET. Phi-4 exhibits the highest improvement on all metrics when reasoning is used. Together with Llama 3.3, they are the highest scoring open-weight models, though Llama 3.3 draws only modest benefits from reasoning. The DeepSeek-R1 models present a mixed picture: only the largest model (32B) improves its BLEU score, though not the other scores.
Models with weaker baselines, i.e., DeepSeek-R1 8B and 14B, Llama 3.1 and 3.2, and Mistral decrease across all metrics when prompted to reason. The scores of NLLB-200 place it midway, behind GPT, Llama 3.3 and Phi, but ahead of DeepSeek-R1 and Mistral. A possible explanation for the scores of DeepSeek-R1 is that, being a reasoning model, the additional reasoning instructions in the user prompt (beyond “please reason step by step”) have a detrimental effect, together with the system prompt, especially for the smaller models.
@@
On the lexical choice benchmark, the ranking of the LLMs is similar to the one on anaphora (see Table 5). Within the GPT family, reasoning improves GPT-4, GPT-4o, and GPT-3.5-turbo across most metrics. While GPT-4-turbo shows an unexpected small decline, GPT-4 seems to make a slightly more effective use of the reasoning instructions, despite GPT-4o’s stronger baseline score. Among open-weight models, Phi-4 again improves on all metrics with the CoT prompt, while LLama 3.3 and DeepSeek-R1 32B improve only in BLEU and chrF, Mistral gains slightly, and the other models degrade, while NLLB-200 is close to the bottom of the ranking. Lexical choice may pose a more difficult challenge compared to pronominal anaphora, as acceptable translations of ambiguous words are more numerous. This may explain why improvements through reasoning are slightly less consistent for the lexical choice task.
6.3.  The “Wise Get Wiser” Effect
In both the anaphora and lexical choice experiments, we observe a surprising effect. Unlike many techniques which tend to improve weaker models but do not benefit the top-scoring ones, here reasoning is more beneficial to the LLMs which already score highest without it. Table 6 shows the Pearson correlations and the Spearman rank correlations between baseline scores without reasoning and the improvements obtained with reasoning ($\Delta$), for both tasks. For anaphora, high coefficients for all metrics (0.59–0.81) confirm a consistent “wise get wiser” pattern: models that start from higher baselines tend to benefit more from reasoning. The effect is weaker for lexical choice, although correlations remain positive (0.21–0.52).
BLEU chrF BERT COMET
Score
Anaphora P 0.81 0.76 0.60 0.59
S 0.68 0.72 0.64 0.74
Lexical P 0.40 0.21 0.31 0.30
choice S 0.52 0.27 0.30 0.25
Table 6: Pearson (P) and Spearman (s) correlations between the baseline scores without reasoning and the improvements due to reasoning ($\Delta$). The large correlations show the “wise get wiser” effect.
Moreover, we confirm that the variations of translation scores between the prompts without reasoning and those with reasoning are consistent across all metrics. This is important because scoring LLM translations uses MT metrics which correlate imperfectly with human preferences, unlike the accuracy metric of the first task. Figure 3, upper part, shows correlations among the $\Delta$ values of all systems, between all pairs of metrics. The rather large coefficients (Pearson: 0.58–0.91, Spearman: 0.66–0.92) indicate that models scoring well with respect to the others on one metric tend to score well on the other metrics as well. Similar values are observed for the lexical choice task, shown in Figure 3, lower part (Pearson: 0.71–0.90, Spearman: 0.71.–0.83).
Figure 3: Correlations between the $\Delta$ values (reasoning minus no-reasoning) obtained with several MT metrics for 12 LLMs, on anaphora translation (upper part) and lexical choice translation (lower part). Each cell shows Pearson (upper triangle) and Spearman (lower triangle) correlations.